# DiffHO: Diffusion-Based Synthesis of Hand-Object Interactions from Textual Descriptions

### v1

@@ 5.6. Ablation Studies
We perform several ablations to justify our technical contributions. To evaluate the controllability gained through guidance, we compare our final method with several variants in Tab. 3. We use a base model which indicates a single stage model without guidance. We then gradually add our contributions, first the grasping and interaction task separation (2-ST), the grasp guidance (GG) and the sub-sequence guidance in the interaction phase (SS) as described in Sec. 4.2. As expected, the grasp reference is being ignored without guidance (top 2 rows in Tab. 3), leading to high grasp errors and low hand correctness. Adding guidance allows for better performance in grasp error and hand accuracy, however, a high indicates a sudden jump of the hands between grasping and interaction. To mitigate this, we introduce the smooth transition technique that leads to the best overall score at the cost of a slight increase in grasp error and decrease in hand correctness. The interpenetration volume is lowest with our final model, indicating the best generalization to unseen objects. To validate our canonicalized representation, we compare against a variant inspired by D-Grasp (Christen et al., 2022), our representation without SDF, using object-relative positions at every step, and finally our representation of object-relative with regards to the initial frame. Our final representation achieves less interpenetration and better contact modeling.
@@
| Pose Representation | Metrics | ||
| IV [] () | ID | CR | |
| D-Grasp (Christen et al., 2022) | 5.56 | 8.60 | 0.055 |
| Ours w/o SDF | 10.79 | 11.42 | 0.073 |
| Ours w/ object-relative pose | 9.38 | 11.41 | 0.077 |
| Ours w/ first frame relative pose (Ours) | 9.03 | 11.39 | 0.086 |
@@ 5.7. Controllability via Text Prompts
In this experiment, we train DiffHO once with our detailed text labels and once with simple text prompts generated from the GRAB objects and action labels. As shown in Tab. 5, our textual descriptions offer significant advantage by enabling control over handedness. To gauge controllability, we measure whether the active hand in the generated motion corresponds to prompted hand. Our model, trained with detailed text annotations achieves a correct overall handedness accuracy of 86.5% against 59.3% accuracy when trained with simple text prompts. We see a performance drop when training on detailed texts and testing on simple texts (0.731) and vice-versa, which was similarly observed in (Shvetsova et al., 2023) due to a style gap. Similarly, we measure action correctness by evaluating the match ratio between the prompted action and the action demonstrated in the generated motion. Using detailed textual descriptions enables our model to generate motions that align more accurately with the text input, demonstrated by a 93.5% accuracy compared to the 54.8% accuracy achieved by our model tested with simple text descriptions. Since the test inputs are much more diverse with the detailed textual descriptions, this furthermore demonstrates the robustness to unseen sentences. We also report the cosine similarity between the different sets of texts as a reference.
| Train | Test | Action | Hand correctness | Cosine | |||||
| Input | Input | Correctness | Right | Left | Both | Total | Similarity | ||
@@
5.8. Motion Generation from Estimated Pose
We run an image-based pose estimator (Pavlakos et al., 2023) on single images from HO3D (Hampali et al., 2020) to obtain object-relative hand poses. We then use these hand poses as reference grasps and pass them through our diffusion framework together with multiple different textual descriptions. This highlights the practicality of our framework-given a single grasp reference, we can generate multiple diverse sequences.
@@ 6. Discussion and Conclusion
In this work, we have introduced a framework to generate plausible hand-object interactions from textual descriptions. Specifically, we have proposed a two-stage diffusion method that separates sequences into a grasping and interaction stage and uses single keyframe grasps as guidance to the diffusion model to improve generalization across unseen objects. Through our compact representation, hands are coupled to object motions resulting in realistic interactions. While we provide a first step towards object generalization, there are limitations of our framework. To remove physical artifacts, physics could be explicitly integrated into the diffusion process (Yuan et al., 2023). The inference time is still relatively slow in diffusion models. To increase efficiency, operating in latent space is an interesting direction to explore (Chen et al., 2023a). Lastly, grasp references for guidance which are far away from the training distribution tend to be ignored by our model. Therefore, exploring new ways to guide the diffusion model towards out-of-distribution samples is a crucial future research direction.
Appendix - Supplementary Material
@@ Appendix A Implementation Details
DiffHO architecture is based on UNet with Adaptive Group Normalization (AdaGN) which was originally proposed in (Dhariwal and Nichol, 2021) and also adapted in (Karunratanakul et al., 2023) for sequence prediction tasks. We adapt this network architecture for our two-stage design, that involves pre-grasp and post-grasp diffusion models. We provide the hyperparameter settings used in our architecture in Table 6.
@@ Appendix B Experimental Details
B.1. Training the action classifier.
We train a standard RNN action recognition classifier on the GRAB dataset and use the final layer of the classifier as the motion feature extractor for calculating action recognition accuracies as well as diversity (DIV) and multimodality (MM) scores, following (Ghosh et al., 2023) and (Guo et al., 2020). We employ the online available code from IMoS and use the same network architecture and parameters to train our model.
B.2. Unseen object split
To assess our model’s ability to generalize to unseen classes, we also introduce a new split comprising unseen objects that were not present during training. Our test set is composed of the following objects: “apple, mug, train, elephant, alarm clock, small pyramid, medium cylinder and large torus”. Note that the training set includes pyramids, cylinders, and torus in other sizes. We exclude small pyramids, medium cylinders, and large torus from the training set to test our models’ ability to generalize to different sizes of the same objects
B.3. Details of the Action Recognition Evaluation Metrics.
Recognition Accuracy.
Given a pretrained classifier, we input motions generated by DiffHO as well as our baselines and report top-1 percentage accuracy. The recognition accuracy indicates the correlation of the action type and the motion.
Diversity.
@@ B.2. Unseen object split
Multimodality.
Different from the diversity metric, multimodality computes variance only within a specific action. An overall score is attained based on averaging the variances across all action types.
@@
FID is computed based on the (Fréchet) distance of the features of real ground-truth motions and generated motions. It has the assumption that the features computed from the real data and synthesized data have a Gaussian distribution. Earlier work demonstrated that for small sample sizes, FID metric is biased and not stable.
@@
| Pre-Grasp | Post-grasp | |
@@
B.4. Visualization Pipeline.
We use AITViewer (Kaufmann et al., 2022) to visualize motions synthesized by DiffHOas well as our baselines. IMoS synthesizes 15 frames of human-object interactions. They determine the hand which is in contact with the object according to GT, which can lead to a jump of the object from one hand to the other for certain actions. To make the frame rates comparable with DiffHO, we upsample the synthesis results of IMoS to 75 frames using spherical interpolation. Therefore, this jump may result in an artifact in which the object is floating from one hand to another.
@@ C.1. Action Recognition Experiments
IMoS (Ghosh et al., 2023) reports results using the whole dataset to train the action classifier on ground-truth motions and testing on the synthesized actions. Since their model initializes the sequences with ground-truth from the test data, this could positively bias their scores. Therefore, we omit test data from training, in our analysis in Table 1 of the main paper, both for our approach and IMoS. We show however in this supplemental material (Table 7) that our model is still able to outperform IMoS even when all the GRAB data, including test data, is included in training the classifier. This is a setup which is more favorable for IMoS as the first frame of synthesized IMoS motion comes from test data. This re-confirms the ability of our method to encode diverse motion features that represents realistic hand-object interactions.
| Method |
DIV () | MM () | KID () | FID () | |||||
Real Mocap | 1.0 | 1.0 | 1.13580.0164 | 0.31050.0163 | - | - | ||||
| IMoS (Ghosh et al., 2023) | 0.7017 | 0.6754 | 1.10530.0128 | 0.28820.0114 | 0.005811 | 0.6267 | |||||
| DiffHO | 0.8125 | 0.9028 | 1.14400.0115 | 0.31760.0127 | 0.005346 | 0.8342 | |||||
Real Mocap | 1.0 | 1.0 | 1.13240.0119 | 0.23550.0044 | - | - | ||||
| IMoS* (Ghosh et al., 2023) | 0.7161 | 0.8000 | 1.11730.0112 | 0.25130.0104 | 0.006038 | 0.7101 | |||||
| DiffHO | 0.7836 | 0.8125 | 1.12250.0120 | 0.21740.0074 | 0.005494 | 0.8821 |
@@
| Method | DIV () | MM () | KID () | FID () | ||||
Ground-truth | 1.06320.0131 | 0.21550.0091 | - | - | |||
| IMoS (Ghosh et al., 2023) | 1.02370.0158 | 0.25380.0103 | 0.008065 | 0.6267 | ||||
| DiffHO | 1.07570.0129 | 0.30370.0079 | 0.006697 | 0.8342 | ||||
Ground-truth | 1.07570.0129 | 0.20020.0063 | - | - | |||
| IMoS* (Ghosh et al., 2023) | 1.04710.0129 | 0.22020.0093 | 0.008641 | 0.6593 | ||||
| DiffHO | 1.09420.0125 | 0.23070.0060 | 0.007503 | 0.8425 |
@@
Given an initial pose of the grasp moment, IMoS only synthesizes post-grasp sequences. To compare our results to IMoS in Table 1 of the main paper, we evaluated our models only on the post-grasp sequence part of the synthesis. In Table 9, we also provide diversity, multimodality, FID, KID and action recognition scores computed over the full sequence including pre- and post-grasp stages. We compare our results against real motions from ground-truth and demonstrate that we can achieve scores similar to ground-truth, which demonstrates our capability to synthesize realistic hand-object interactions.
@@
| DiffHO (Full-sequence) | 0.7720 | 0.8031 | 1.09480.0120 | 0.15550.0057 | 0.0067 | 0.7993 |
C.2. Text Input Evaluation Experiments
In Table 10, we provide further details for the text input to the diffusion model. We first compare IMoS and our model trained on the subject split and tested on the object split. Notably, our model outperforms IMOS by 9.2% in action correctness, a significant improvement. When trained and tested on our unseen object splits, we find that our model remains competitive with IMoS. However, we notice that our model struggles with classes like apple and train, where new motions are introduced that differ from those in the training set. Our model must learn to generalize from eating bananas to apples, and while it achieves this, its performance decreases to 81.8%. Similarly, for the train class, our model’s accuracy drops to 50% since the unique motions of playing with a train were not observed during training. Nevertheless, our model demonstrates the ability to generalize to different object sizes, as it generates motions corresponding to recognizable actions for varied sizes of cylinders, pyramids, and torus in the training set.
Finally, we train and test our model with the augmented sentences we proposed for GRAB, leading to improved generalization, especially for the classes “train” and “apple”. These new sentences provide detailed motion descriptions, such as “The person picks up the train using their right hand, then rides it through the air close to the table surface, and finally puts it on the table using their right hand.” The augmented sentences enhance generalization and provide control over specifics like handedness, motion type, and physical location. Testing a model trained solely with baseline sentences on augmented sentences reveals a decrease in generalization, as anticipated. We observed a significant 38.7% performance gap between baseline and augmented sentences, showing the importance of text style. This highlights the limitations of relying solely on simple sentences for training and underscores the value of incorporating diverse and detailed descriptions to enhance model performance and generalization capabilities.
C.3. Detailed Investigation User Study
In our user study, we provide each user with 30 side-by-side comparisons. In our investigation, our model showed low realism but high diversity for the “pass” action, influenced by users perceiving “pass” action label in the GRAB dataset as transferring an object between hands, not to another person. In the “offhand” category, our model excelled in realism, while IMoS exhibited artifacts. This is due to the fact that IMoS determines the hand which is in contact with the object according to ground-truth, which leads to a jump of the object for the “off-hand” category.
C.4. More Qualitative Results and Failure Cases
We provide more qualitative results of our model with simple instructions in Fig. 5(a), for grasp guidance in Fig. 5(b), and with detailed instructions in Fig. 5(c). Furthermore, we present three failure cases: the generated action does not match the text prompt (Fig. 6(a)), the grasp reference is ignored during guidance (Fig. 6(b)), or the hand in the text prompt is ignored (Fig. 6(c)).
| Split | Textual input form |
Action correctness over object classes | ||||||||||||||||||
| Method | Train split | Test split | Train input | Test input | Overall | apple | mug | train |
elephant |
| IMOS* | subject | object | simple | simple | 90.3 | 90.9 | 89.5 | 77.8 | 100 | 80.0 | 100 | 83.3 | 100 | ||||||||
| Ours | subject | object | simple | simple | 99.5 | 100 | 97.4 | 100 | 100 | 100 | 100 | 100 | 100 | ||||||||
| Ours | object | object | simple | simple | 89.7 | 81.8 | 92.1 | 50.0 | 100 | 100 | 100 | 100 | 100 | ||||||||
| Ours | object | object | simple | comprehensive | 54.8 | 40.9 | 55.3 | 16.7 | 52.6 | 80.0 | 75 | 50.0 | 68.8 | ||||||||
| Ours | object | object | comprehensive | comprehensive | 93.5 | 95.5 | 94.7 | 77.8 | 100 | 95.0 | 93.6 | 100 | 93.6 | ||||||||

### v2

@@ 5.6. Ablation Studies
We perform several ablations to justify our technical contributions. To evaluate the controllability gained through guidance, we compare our final method with several variants in Tab. 3 for the full setting that generates the whole sequence with grasping and interaction (Setting 2). We use a base model which indicates a single stage model without guidance. We then gradually add our model components: (i) 2-stage design that decouples grasping and interaction (2-ST), (ii) grasp guidance (GG), and, (iii) subsequence imputing in the interaction phase (SI) as described in Sec. 4.2. As expected, the grasp reference is being ignored without guidance (top 2 rows in Tab. 3), leading to high grasp errors and low hand correctness. Adding guidance allows for better performance in grasp error and hand accuracy, however, a high indicates a sudden jump of the hands between grasping and interaction. Our subsequence imputing mitigates this and leads to the best overall score at the cost of a slight increase in grasp error and decrease in hand correctness. The IV is lowest with our final model, indicating the best generalization to unseen objects. To validate our representation, we further compare our approach, which normalizes hands with respect to the object position at first frame, against a number of baselines: (i) a variant inspired by D-Grasp (Christen et al., 2022) that involves a redundant representation involving joint angles in Euler space and angular velocities, (ii) our representation without SDF, and, (iii) our representation with per-frame object relative poses. As shown in Tab. 4, our final representation achieves the best overall scores and trade-off with low interpenetration and high contact ratios. As discussed in Sec. 4.1, our representation relative to initial frame is less variant with specific object shapes and more correlated with motion involved in hand-object interaction compared to per-frame object-relative representation. This effectively improves generalization for HOI across different objects. In addition, SDF feature further improves accuracy by reasoning about local shape.
The D-Grasp representation has the lowest interpenetration scores, but as indicated by the contact ratio and our qualitative results in the video, the hands are often disconnected from the objects, potentially due its challenge in optimizing in high-dimensional, redundant feature space. Additionally, our PCA space helps reduce unnatural hand poses that occur when predicting joints in angle space directly (see supplemental video). While the PCA space is beneficial in data-sparse settings, it may hinder more fine-grained finger manipulations compared to joint angle predictions.
@@
5.7. Generation from Estimated and Pre-Synthesized Pose
We show two qualitative applications of DiffH2O in video. First, we run an image-based pose estimator (Pavlakos et al., 2024) on single images from HO3D (Hampali et al., 2020) to obtain object-relative hand poses. Second, we generate grasping poses using GraspXL (Zhang et al., 2024a). We then pass these hand poses through our method as reference grasps in grasp guidance (see Sec. 4.3), together with multiple different textual descriptions. This highlights the practicality of our framework given a single grasp reference or grasp motion, to generate multiple diverse sequences. See supplementary material for details about how we obtain grasp references and further discussions on their benefits and limitations.
@@ 5.8. Controllability via Text Prompts
In this experiment, we train DiffH2O once with our detailed text descriptions and once with simple text prompts generated from the GRAB objects and action labels. We measure whether the active hand in the generated motion corresponds to the prompted hand to gauge controllability. Results in Tab. 5 show that detailed text descriptions significantly improve control over handedness, achieving an accuracy of vs. with simple prompts. Performance drop when training on simple texts and testing on detailed texts (0.702) and vice-versa, which was similarly observed in (Shvetsova et al., 2023) due to a style gap. We also measure action correctness by evaluating action recognition accuracy (AR, see Sec. 5.3). Using detailed descriptions enables our model to generate motions that align more accurately with the text input, demonstrated by a 0.887 action accuracy compared to 0.516 with our model tested with simple prompts. Since the test inputs are much more diverse with the detailed textual descriptions, this furthermore demonstrates the robustness to unseen sentences. We also report the cosine similarity between the different sets of texts as a reference.
| Pose Representation (unseen objects) | Metrics | ||
| CR | IV [] | ID | |
| D-Grasp (Christen et al., 2022) | 0.055 | 5.56 | 8.60 |
| Ours w/o SDF | 0.073 | 10.79 | 11.42 |
| Ours w/ object-relative pose | 0.077 | 9.38 | 11.41 |
| Ours w/ first frame relative pos. (Ours) | 0.086 | 9.03 | 11.39 |
| Train | Test | AR | Hand correctness | Cosine | |||||
| Input | Input | Right | Left | Both | Total | Similarity | |||
@@ 6. Discussion and Conclusion
We have introduced a framework to generate plausible hand-object interactions from textual descriptions. Specifically, we have proposed a two-stage diffusion framework that decouples sequences into grasping and interaction, and can use single grasps as guidance to the diffusion model to improve generalization to unseen objects. While we provide a first step towards HOI synthesis generalization, there are limitations of our framework. To remove physical artifacts, physics could be explicitly integrated into the diffusion process (Yuan et al., 2023). The inference time is still relatively slow. To increase efficiency, operating in latent space is an interesting direction to explore (Chen et al., 2023a).
@@ Appendix A Implementation Details
While we experiment with a transformer backbone as proposed in MDM (Tevet et al., 2023), DiffH2O’s main architecture is based on UNet with Adaptive Group Normalization (AdaGN) which was originally proposed in (Dhariwal and Nichol, 2021) and also adapted in (Karunratanakul et al., 2023) for sequence prediction tasks. We adapt this network architecture for our two-stage design, that involves grasping and interaction diffusion models. We provide the hyperparameter settings used in our architecture in Table 6.
@@ Appendix B Experimental Details
B.1. Obtaining Grasp References
As mentioned in the main part of the paper, we show three different ways grasp references can be obtained, namely via existing datasets, pose estimation, or grasp synthesis. Note that these procedures for acquiring hand pose references have already been established in related works (Christen et al., 2022; Xu et al., 2023b), albeit for the sole purpose of grasping and not interaction.
Hand-Object Datasets
The most straightforward way to attain grasp references is to make use of existing hand-object datasets (see Sec. 2). There are some single frame datasets like ObMan (Hasson et al., 2019) or AffordPose (Ye et al., 2023b) that could be directly used as grasp references. However, most existing datasets contain sequential data, hence, we need to identify which frames are potential candidates for hand-object grasps. In our paper, we run guidance experiments on our test split of the GRAB dataset (Taheri et al., 2020). To acquire suitable frames for guidance, we apply the heuristics described in Sec. 4.2.1. In this case, we can directly use the preprocessed data from GRAB as single keyframes for guidance.
Image-Based Pose Estimation
For this experiment, we need to extract the grasp references from images. We run an image-based pose estimator (Pavlakos et al., 2024) on single images from HO3D (Hampali et al., 2020) to obtain object-relative hand poses. In this experiment, we focus on extracting a single hand pose from the image. Please note that we assume the object mesh to be known and at a normalized position to be suitable for our experiments. Given the object-relative hand pose, we need to make three modifications to convert it to the same input space our model has. First, we have to compute the features for our approximated signed-distance space by finding the closest vertex on the object mesh for each joint position. Second, we convert the hand pose from joint angle space to MANO’s PCA space. Lastly, because the GRAB dataset is normalized to approach objects from one direction, we have to adjust the orientation of the hand and object to roughly be in line with this direction. This could be alleviated by adding data augmentation to the training data in the future. We also leave exploring object mesh reconstruction for future work, for example, by utilizing methods that reconstruct the object mesh along with the hand pose estimation (Fan et al., 2024).
Grasp Synthesis
In this experiment, we use a grasp synthesis method, GraspXL (Zhang et al., 2024a), to obtain grasp references for guidance. To this end, we run GraspXL on several objects to obtain grasping motions. As we only require one frame for the guidance, we take the last frame of the grasping motion as the keyframe. Since GraspXL allows conditioning on approaching direction, we select grasps whose approaching direction aligns with the GRAB dataset. Again, we compute the approximated signed distance by finding the closest vertex on the object mesh for each joint position and convert the hand pose from joint angle to MANO PCA space. There is no further conversion needed, as the object poses are already normalized. Similar to the experiment with hand poses from images, we only acquire an object-relative grasp and its respective features for one hand.
B.2. Additional Motion Feature Evaluation Metrics.
For completion, we also provide metrics based on an action recognition classifier following previous motion synthesis works (Ghosh et al., 2023; Tevet et al., 2023).
Action Recognition Accuracy (AR).
Given a pretrained classifier, we input motions generated by DiffH2O as well as our baselines and report top-1 percentage accuracy. The recognition accuracy indicates the correlation of the action type and the motion.
Diversity (DIV).
@@ B.1. Obtaining Grasp References
Multimodality (MM).
Different from diversity, multimodality computes variance only within a specific action. An overall score is attained based on averaging the variances across all action types.
@@
FID is computed based on the (Fréchet) distance of the features of real ground-truth motions and generated motions. It has the assumption that the features computed from the real data and synthesized data have a Gaussian distribution. Earlier work demonstrated that for small sample sizes, FID metric is biased and not stable (Noguchi and Harada, 2019).
@@
B.3. Metrics - Training the Action Classifier.
To provide motion feature metrics (see Sec. B.2) based on an action classifier, we train a standard RNN action recognition classifier on the GRAB dataset. We use the final layer of the classifier as the motion feature extractor for calculating action recognition accuracies as well as diversity (DIV) and multimodality (MM) scores, following (Ghosh et al., 2023) and (Guo et al., 2020). We employ the online available code from IMoS and use the same network architecture and parameters to train our model.
| Grasping | Interaction | |
@@
B.4. Annotation Examples
We provide a few examples of our textual annotations in Table 10. Please see our code release for the entire set of annotations***The code and data will be available here: https://diffh2o.github.io.
B.5. Visualization Pipeline.
We use AITViewer (Kaufmann et al., 2022) to visualize motions synthesized by DiffH2O as well as our baselines. IMoS synthesizes 15 frames of human-object interactions. They determine the hand which is in contact with the object according to GT, which can lead to a jump of the object from one hand to the other for certain actions. To make the frame rates comparable with DiffH2O, we upsample the synthesis results of IMoS to 75 frames using spherical interpolation. Therefore, this jump may result in an artifact in which the object is floating from one hand to another.
@@ C.1. Action Recognition Experiments
IMoS (Ghosh et al., 2023) reports results using the whole dataset to train the action classifier on ground-truth motions and testing on the synthesized actions. Since their model initializes the sequences with ground-truth from the test data, this could positively bias their scores. Therefore, we omit test data from training, in our analysis in Table 1 of the main paper, both for our approach and IMoS. We show however in this supplemental material (Table 7) that our model is still able to outperform IMoS even when all the GRAB data, including test data, is included in training the classifier. This is a setup which is more favorable for IMoS as the first frames of synthesized IMoS motion is taken from the test data. This re-confirms the ability of our method to encode diverse motion features that represents realistic hand-object interactions. Furthermore, we train two models, one that uses only hand 3D joints as input and another incorporating object 3D position as additional input. We demonstrate that DiffH2O’s synthesized object motions outperform IMoS in both settings. This confirms the benefits of jointly modeling hand-object motion instead of decoupling it into separate components. Note also that as expected, the classifier can correctly identify all test sequences of the real mocap data, since it has seen the test data during training in the IMoS setting.
@@
| Method |
DIV () | MM () | KID () | FID () | |||
| Unseen subject split | Real Mocap | 1.0 | 1.0 | 1.13580.0164 | 0.31050.0163 | - | - | ||
| IMoS (Ghosh et al., 2023) | 0.7017 | 0.6754 | 1.10530.0128 | 0.28820.0114 | 0.005811 | 0.6267 | |||
| DiffH2O | 0.8125 | 0.9028 | 1.14400.0115 | 0.31760.0127 | 0.005346 | 0.8342 | |||
| Unseen object split | Real Mocap | 1.0 | 1.0 | 1.13240.0119 | 0.23550.0044 | - | - | ||
| IMoS* (Ghosh et al., 2023) | 0.7161 | 0.8000 | 1.11730.0112 | 0.25130.0104 | 0.006038 | 0.7101 | |||
| DiffH2O | 0.7836 | 0.8125 | 1.12250.0120 | 0.21740.0074 | 0.005494 | 0.8821 |
| Method | DIV () | MM () | KID () | FID () | |
| Unseen subject split | Real Mocap | 1.06320.0131 | 0.21550.0091 | - | - |
| IMoS (Ghosh et al., 2023) | 1.02370.0158 | 0.25380.0103 | 0.008065 | 0.6267 | |
| DiffH2O | 1.07570.0129 | 0.30370.0079 | 0.006697 | 0.8342 | |
| Unseen object split | Real Mocap | 1.07570.0129 | 0.20020.0063 | - | - |
| IMoS* (Ghosh et al., 2023) | 1.04710.0129 | 0.22020.0093 | 0.008641 | 0.6593 | |
| DiffH2O | 1.09420.0125 | 0.23070.0060 | 0.007503 | 0.8425 |
@@
Given an initial pose of the grasp moment, IMoS only synthesizes interaction sequences. To compare our results to IMoS in Table 1 of the main paper, we evaluated our models only on the interaction sequence part of the synthesis. In Table 9, we also provide diversity, multimodality, FID, KID and action recognition scores computed over the full sequence including grasp and interaction stages. We compare our results against motions from motion capture and observe that we can achieve scores similar to real mocap, which demonstrates our capability to synthesize realistic hand-object interactions.
@@
| DiffH2O (Full-sequence) | 0.7720 | 0.8031 | 1.09480.0120 | 0.15550.0057 | 0.0067 | 0.7993 |
| Object Name | Annotation |
| camera | The person uses their left hand to pick up the camera, brings it closer to their eyes to take a photo, and then places it back on the table with their right hand. |
| stapler | The person repeatedly picks up and places the stapler on the table always with their right hand. |
| mouse | The person picks up the mouse with their right hand, passes it to someone on their right below chest level, and then places it back on the table with their right hand. |
| alarm clock | The person repeatedly picks up the alarm clock with their right hand, lifts it, and places it back on the table with their right hand. |
| duck | Using the right hand, the person picks up the duck, examines it closely, and then places the duck back down on the table. |
| bowl | The person picks up the bowl with both hands, drinks from it, and places it back on the table with both hands. |
| headphones | The person picks up the headphones with their left hand, puts them on with both hands, and places them on the table with both hands. |
| banana | The person picks up the banana with their right hand, passes it to their left hand, then hands it to someone on their left side at shoulder level with their left hand, and finally places it back on the table with their right hand. |
C.2. Analysis on Generalizability
When we analyze per-unseen-object performance on GRAB, our method effectively generalizes to unseen objects with various sizes and shapes (e.g., apple, clock, train, elephant, mug, etc.). We also evaluated the influence of object scale on three test objects by using objects of the same shape but different scales during training: our model performs well on medium cylinder and large torus (584cm³), but faces challenges on small pyramid (64cm³, video: 04.40, low contact-ratio). The average bounding-box volumes for our train and test sets are 741cm³ and 625cm³, respectively. In conclusion, our model adapts well to unseen shapes and a wide range of sizes, but largely out-of-distribution sizes (e.g., also large_cracker_box from HO3D: 2208cm³) lead to less favorable physics metrics.
C.3. Discussion on Grasp References
Here we provide a more detailed discussion about the role of grasp references that are used for guidance.
Strengths:
Grasp guidance is a technique to increase the controllability and improve generalization on unseen objects if grasp references are available. Note, however, that our model without guidance already significantly outperforms baselines (cf. Table 1 in the main paper), making it an ideal tool for potential downstream use cases. Grasp guidance yields better generalization (e.g., less interpenetration as shown in Table 2 in the main paper), because it provides a signal to the diffusion model about how an object should be grasped. Especially for unseen object geometry, this can serve as important information about feasible grasping poses of the hand. One of the other main advantages of grasp guidance is that it allows defining the grasping hand(s) on the object, which may be important if rich textual annotations such as the ones presented in Sec. 4.4 are not available.
Weaknesses and Failures:
Grasp guidance faces an important trade-off: while the signal can be helpful for object generalization and controllability of handedness, it requires access to single hand-object poses. We detail in Section B.1 how such grasp references can be obtained. Hence, there are conversion steps that are necessary to make guidance work on unseen object geometry, which require some manual effort. For instance, due to the small dataset size of GRAB and the similar approaching direction for all objects, an alignment step is necessary to make the guidance signal useful. In the future, larger dataset distributions or additional data augmentation techniques may help alleviate this. Furthermore, as our model is conditioned on both the text description and the grasp guidance, one needs to make sure that the two align: for example, if the grasp reference shows grasping a cup from above, where the prompted action is ”drink from the cup”, the model is forced to select which signal to follow. In practice, as the guidance is only a weak guidance signal at a single frame, the model will favor the text description over the guidance. Another limiting factor is the size of the unseen objects as discussed in Section C.2. For example, if the object is too large, the target grasp may be reached in the grasping phase, but after imputing the grasping into the interaction stage, the model may neglect the out-of-distribution grasp and output a grasp that displays physical artifacts such as interpenetration.
Overall, grasp guidance is a technique that shows the potential to provide a link between how an object should be grasped and subsequent actions that can be prompted via text. While there are limiting factors in its current state, i.e., manual alignment and matching text prompt with the grasp reference, future work could focus on enhancing existing datasets and using improved guidance techniques (Karunratanakul et al., 2024) to further increase performance.